# Functional Stochastic Localization

### v1

@@ 1 Introduction
This line of inquiry is motivated by the rich literature on non-Euclidean optimization, which has had wide-ranging applications such as sparse recovery [CRT06], matrix completion [ANW10], combinatorial optimization [KLO+14, KPS+19], multi-armed bandit problems [BC12], and more. Many of these applications are intimately related to the framework of mirror descent, a “functional” generalization of gradient descent induced by a non-Euclidean regularizer $\psi:\mathbb{R}^{d}\to\mathbb{R}$. A parallel question would then be: is there a functional generalization of stochastic localization, e.g., a distribution over linear tilts $\mathbf{c}_{\tau}\in\mathbb{R}^{d}$ that satisfies, for some convex $\psi:\mathbb{R}^{d}\to\mathbb{R}$, and $\pi$ a density on $\mathbb{R}^{d}$,
$\mathbb{E}\left[\pi_{\tau}\right]=\pi,\text{ pointwise on }\mathbb{R}^{d},\text{ where }\pi_{\tau}:=\mathcal{T}_{\mathbf{c}_{\tau}}\exp{\left(-\tau\psi\right)}\pi?$ (1)
@@ 1.1 Our approach
Our framework is motivated by and related to [GLL+25], but substantially broadens its flexibility. To better explain this relationship, it is helpful to outline a typical mixing-time result for Euclidean proximal samplers. Let $\pi$ be a density on $\mathbb{R}^{d}$ that satisfies an $\alpha$-Poincaré inequality (PI, Definition 3). The proximal sampler of [LST21] can be viewed as performing Gibbs sampling on an extended space, defined by a random tilt $\mathbf{c}_{\tau}$ at time $\tau>0$, and a particle $\mathbf{x}\in\mathbb{R}^{d}$. The mixing time of the proximal sampler (in $\chi^{2}$ divergence) then scales $\propto\frac{\tau}{\alpha}$ [LST21, CCS+22, CE22]. This leads to a tradeoff: larger $\tau$ leads to slower mixing, but induces additional, helpful, Gaussian regularization in the tilted densities $\pi_{\tau}$. Choosing $\tau$ to balance the cost of Gibbs sampling with the mixing time then enables a host of applications.
@@
Our framework allows for taking arbitrarily large $\tau$, with a twist: although the underlying space $\mathbb{R}^{d}$ is continuous, $\tau\in\mathbb{N}$ in our process now indexes a discrete time. We then construct the process defining $\mathbf{y}_{\tau}$ in (1) using $\tau$ convolved draws from an exponential family (Algorithm 1, see also the alternate definition of our process in (8)). This construction is motivated by an information-theoretic viewpoint of stochastic localization as posterior sampling, given progressively denoised observations from a Gaussian channel [AM22] (see also [STZ25], Section 3). Our key observation is that, unlike the SDE-based definitions of stochastic localization, the perspective based on posterior sampling extends cleanly to discrete time, where $\tau\in\mathbb{N}$ is the number of observations. We formalize these ideas in Section 3 by leveraging properties of exponential families, where we prove that Algorithm 1 meets the requirement (1) and is a localization process in the sense of [CE22].
@@ 4 Gibbs Sampling
In this section, we revisit general two-state Gibbs sampling, and provide several structural facts relating the variance contraction of the “primal process” that samples from one of two state spaces, to the variance contraction of the corresponding “dual process” on the other state space. These facts will be used to analyze Algorithm 1 in Section 5.
@@ 4.1 Variance contraction via spectral gap
Now, we interlude with a discussion on the mixing rate for Markov chains associated to general localization processes by completing the argument in Proposition 19, [CE22]. That is, Proposition 19, [CE22] gives a generic way of converting a variance conservation bound in a localization scheme to a spectral gap on the associated Markov chain. We provide a more explicit variant of this calculation that in fact holds for generic Markov chains, not just Gibbs samplers.
@@
We apply Lemma 6 to prove a convergence rate on Gibbs sampling from the LLT-induced joint density (7) in Section 5. Interestingly, even the specialization of our result to Euclidean stochastic localization yields a rate that is twice as fast as derived in existing applications of localization schemes machinery (e.g., Theorem 58, [CE22]), which owes to our sharper characterization in Lemma 6.
@@ 4 $\mathbf{x}_{k+1}\sim\mathcal{T}_{\mathbf{y}_{k}}\exp(-\tau\psi)\pi$
From the construction given in (8), we draw $\mathbf{w}\sim\pi$ and $\mathbf{c}\mid\mathbf{w}\sim\tilde{\pi}_{(\tau)}^{Y\mid X=\mathbf{w}}$, and set $\pi_{\tau}=\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{c}}$. Then
@@
$=\mathbb{E}_{(\mathbf{w},\mathbf{c})\sim\tilde{\pi}_{(\tau)}}\left[\frac{\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{c}}(\mathbf{x}^{\prime})\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{c}}(\mathbf{x})}{\pi(\mathbf{x})}\right]$
$=\mathbb{E}_{\mathbf{c}\sim\tilde{\pi}_{(\tau)}^{Y}}\left[\frac{\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{c}}(\mathbf{x}^{\prime})\tilde{\pi}_{(\tau)}(\mathbf{x},\mathbf{c})}{\tilde{\pi}_{(\tau)}^{Y}(\mathbf{c})\pi(\mathbf{x})}\right]$
$=\int\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{c}}(\mathbf{x}^{\prime})\tilde{\pi}_{(\tau)}^{Y\mid X=\mathbf{x}}(\mathbf{c})\,\textup{d}\mathbf{c}.$
@@
$\boldsymbol{\Sigma}(\mathbf{x})$ $:=\mathbf{Cov}_{\mathbf{y}\sim\tilde{\pi}^{Y\mid X=\mathbf{x}}}[\mathbf{y}]=\tau\nabla^{2}\psi(\mathbf{y}),$
@@
This matches the Poincaré constant in the Euclidean case, where existing proofs have realized $\tilde{\pi}^{Y}$ explicitly as a Lipschitz transformation of a Gaussian convolution of $\pi$ [STZ25, Proposition 1] or via multiscale Bakry-Émery [BBD24, Theorem 3.6]. Our proof demonstrates that this constant holds more generally via a linear algebraic proof, using properties of the LLT. In Appendix D we show how alternative approaches that work for Gaussians, e.g., via convolution or convexity, break down in our more general setting.
@@ Appendix B Proof of Theorem 1 Weakening via Conductance
Consider a Markov chain on $\mathbb{R}^{d}$ with transition kernels $\{\mathsf{t}_{\mathbf{x}}\}_{vx\in\mathbb{R}^{d}}$ and stationary distribution $\pi$. Suppose the following two conditions hold for some metric $\textup{m}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$.
@@ Appendix C Entropic Contraction Under Assumption 2
To motivate Assumption 2, we first consider the case $\alpha=1$ and $V=\psi$. Then Assumption 2 simply posits that $\psi^{\sharp}=\varphi^{\sharp\sharp}$ is “less convex” than $\varphi$, in the sense that $\varphi-\varphi^{\sharp\sharp}$ is convex. In other words, taking the LLT of a convex function $\varphi$ twice gives a function which is smaller than $\varphi$ in the convex partial ordering. There are some analogies present in the literature between the Fenchel conjugation operator (for which $\varphi^{**}=\varphi$) and the LLT operator, e.g., [CCS+22], Section B.2 and [GLL+25], Section 3.1. Unfortunately, essentially any non-quadratic example witnesses $\varphi^{\sharp\sharp}\neq\varphi$. [^3] An easy way to see this is to take any $\varphi$ that is not self-concordant. Then $\varphi^{\sharp\sharp}=\varphi$ is ruled out simply by Fact 2.
@@
Now let $\gamma_{k+0.5}$ be the coupling of $(\mu_{k+0.5},\mu^{\prime}_{k+0.5})$ which first draws $(\mathbf{x},\mathbf{x}^{\prime})\sim\gamma_{k}$, and then conditioned on these values, draws and outputs $(\mathbf{y},\mathbf{y}^{\prime})$ from the coupling realizing $W^{\varphi^{*\tau}}(\tilde{\pi}_{(\tau)}^{Y\mid X=\mathbf{x}}\|\tilde{\pi}_{(\tau)})$.
Conditioning on the values of $(\mathbf{y},\mathbf{y}^{\prime})\sim\gamma_{k+0.5}$, we have by Lemma 2, Fact 13, and Fact 12 that
@@ Appendix D Comparison of Lemma 10 with Gaussian Setting
We use the second line of (29) and follow the proof of [STZ25, Proposition 1]. Here, $\tilde{\pi}_{(\tau)}^{Y}$ is almost of the form of a Gaussian convolution. First define
@@
Then this is the convolution of two densities that satisfy Poincaré inequalities with constants $\alpha$ and $\tau$, respectively. Therefore, by Lemma 2, $\nu_{\tau}$ satisfies a $\frac{\alpha\tau}{\alpha+\tau}$-Poincaré inequality. Now $\nu_{\tau}$ and $\tilde{\pi}_{(\tau)}^{Y}$ are equivalent up to a $\frac{1}{\tau}$-Lipschitz mapping; thus, by [CHE23, Proposition 2.3.3], $\tilde{\pi}_{(\tau)}^{Y}$ satisfies a $\frac{\alpha}{\tau(\alpha+\tau)}$-Poincaré inequality. ∎
@@
The proof via convexity is brittle in two respects. Generally, Assumption 2 may not hold. Moreover, it is not necessarily the case that $\nabla^{2}\varphi^{*\tau}\succeq\frac{1}{\tau}\nabla^{2}\varphi$. Indeed, consider the proof of [STZ25, Lemma 25] for general $\varphi$. We would like

### v2

@@ 1 Introduction
This line of inquiry is motivated by the rich literature on non-Euclidean optimization, which has had wide-ranging applications such as sparse recovery [CRT06], matrix completion [ANW10], combinatorial optimization [KLO+14, KPS+19], multi-armed bandit problems [BC12], and more. Many of these applications are intimately related to the framework of mirror descent, a “functional” generalization of gradient descent induced by a non-Euclidean regularizer $\psi:\mathbb{R}^{d}\to\mathbb{R}$. A parallel question would then be: is there a functional generalization of stochastic localization, e.g., a distribution over linear tilts $\mathbf{y}_{\tau}\in\mathbb{R}^{d}$ that satisfies, for some convex $\psi:\mathbb{R}^{d}\to\mathbb{R}$, and $\pi$ a density on $\mathbb{R}^{d}$,
$\mathbb{E}\left[\pi_{\tau}\right]=\pi,\text{ pointwise on }\mathbb{R}^{d},\text{ where }\pi_{\tau}:=\mathcal{T}_{\mathbf{y}_{\tau}}\exp{\left(-\tau\psi\right)}\pi?$ (1)
@@ 1.1 Our approach
Our framework is motivated by and related to [GLL+25], but substantially broadens its flexibility. To better explain this relationship, it is helpful to outline a typical mixing-time result for Euclidean proximal samplers. Let $\pi$ be a density on $\mathbb{R}^{d}$ that satisfies an $\alpha$-Poincaré inequality (PI, Definition 3). The proximal sampler of [LST21] can be viewed as performing Gibbs sampling on an extended space, defined by a random tilt $\mathbf{y}_{\tau}$ at time $\tau>0$, and a particle $\mathbf{x}\in\mathbb{R}^{d}$. The mixing time of the proximal sampler (in $\chi^{2}$ divergence) then scales $\propto\frac{\tau}{\alpha}$ [LST21, CCS+22, CE22]. This leads to a tradeoff: larger $\tau$ leads to slower mixing, but induces additional, helpful, Gaussian regularization in the tilted densities $\pi_{\tau}$. Choosing $\tau$ to balance the cost of Gibbs sampling with the mixing time then enables a host of applications.
@@
Our framework allows for taking arbitrarily large $\tau$, with a twist: although the underlying space $\mathbb{R}^{d}$ is continuous, $\tau\in\mathbb{N}$ in our process now indexes a discrete time. We then construct the process defining $\mathbf{y}_{\tau}$ in (1) using $\tau$ convolved draws from an exponential family (Algorithm 1, see also the alternate definition of our process in (8)). This construction is motivated by an information-theoretic viewpoint of stochastic localization as posterior sampling, given progressively denoised observations from a Gaussian channel [AM22] (see also Section 3, [STZ25]). Our key observation is that, unlike the SDE-based definitions of stochastic localization, the perspective based on posterior sampling extends cleanly to discrete time, where $\tau\in\mathbb{N}$ is the number of observations. We formalize these ideas in Section 3 by leveraging properties of exponential families, where we prove that Algorithm 1 meets the requirement (1) and is a localization process in the sense of [CE22].
@@ 4 Gibbs Sampling
In this section, we revisit general two-state Gibbs sampling, and provide several structural facts relating the variance contraction of the “primal process” that samples from one of two state spaces, to the variance contraction of the corresponding “dual process” on the other state space. These facts will be used to analyze Algorithm 1 in Section 5. We remark much of the content in this section has been derived in the information theory literature through the lens of strong data processing inequalities [RAG16, PW25]; we include brief rederivations to keep our exposition self-contained.
@@ 4.1 Variance contraction via spectral gap
Now, we interlude with a discussion on the mixing rate for Markov chains associated to general localization processes by completing the argument in Proposition 19, [CE22]. That is, Proposition 19, [CE22] gives a generic way of converting a variance conservation bound in a localization scheme to a spectral gap on the associated Markov chain. We provide a more explicit variant of this calculation that in fact holds for generic Markov chains, not just Gibbs samplers. For an analogous derivation in the discrete space setting, see Example 33.8, [PW25].
@@
We apply Lemma 6 to prove a convergence rate on Gibbs sampling from the LLT-induced joint density (7) in Section 5. Interestingly, for variance contraction specifically, the specialization of our result to Euclidean stochastic localization yields a rate that is twice as fast as derived in existing applications of localization schemes machinery (e.g., Theorem 58, [CE22]), which owes to our sharper characterization in Lemma 6. We remark that the same generic “forwards contraction implies backwards contraction” characterization of Gibbs samplers is false in full generality, in the case of entropy instead of variance, cf. [CCG+25] for a counterexample.
@@ 4 $\mathbf{x}_{k+1}\sim\mathcal{T}_{\mathbf{y}_{k}}\exp(-\tau\psi)\pi$
From the construction given in (8), we draw $\mathbf{w}\sim\pi$ and $\mathbf{y}\mid\mathbf{w}\sim\tilde{\pi}_{(\tau)}^{Y\mid X=\mathbf{w}}$, and set $\pi_{\tau}=\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{y}}$. Then
@@
$=\mathbb{E}_{(\mathbf{w},\mathbf{c})\sim\tilde{\pi}_{(\tau)}}\left[\frac{\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{y}}(\mathbf{x}^{\prime})\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{y}}(\mathbf{x})}{\pi(\mathbf{x})}\right]$
$=\mathbb{E}_{\mathbf{c}\sim\tilde{\pi}_{(\tau)}^{Y}}\left[\frac{\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{y}}(\mathbf{x}^{\prime})\tilde{\pi}_{(\tau)}(\mathbf{x},\mathbf{y})}{\tilde{\pi}_{(\tau)}^{Y}(\mathbf{y})\pi(\mathbf{x})}\right]$
$=\int\tilde{\pi}_{(\tau)}^{X\mid Y=\mathbf{y}}(\mathbf{x}^{\prime})\tilde{\pi}_{(\tau)}^{Y\mid X=\mathbf{x}}(\mathbf{y})\,\textup{d}\mathbf{y}.$
@@
$\boldsymbol{\Sigma}(\mathbf{x})$ $:=\mathbf{Cov}_{\mathbf{y}\sim\tilde{\pi}^{Y\mid X=\mathbf{x}}}[\mathbf{y}]=\tau\nabla^{2}\psi(\mathbf{x}),$
@@
This matches the Poincaré constant in the Euclidean case, where existing proofs have realized $\tilde{\pi}^{Y}$ explicitly as a Lipschitz transformation of a Gaussian convolution of $\pi$, Proposition 1, [STZ25] or via multiscale Bakry-Émery, Theorem 3.6, [BBD24]. Our proof demonstrates that this constant holds more generally via a linear algebraic proof, using properties of the LLT. In Appendix D we show how alternative approaches that work for Gaussians, e.g., via convolution or convexity, break down in our more general setting.
@@ Appendix B Proof of Theorem 1 Weakening via Conductance
Consider a Markov chain on $\mathbb{R}^{d}$ with transition kernels $\{\mathsf{t}_{\mathbf{x}}\}_{\mathbf{x}\in\mathbb{R}^{d}}$ and stationary distribution $\pi$. Suppose the following two conditions hold for some metric $\textup{m}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$.
@@ Appendix C Entropic Contraction Under Assumption 2
To motivate Assumption 2, we first consider the case $\alpha=1$ and $V=\psi$. Then Assumption 2 simply posits that $\psi^{\sharp}=\varphi^{\sharp\sharp}$ is “less convex” than $\varphi$, in the sense that $\varphi-\varphi^{\sharp\sharp}$ is convex. In other words, taking the LLT of a convex function $\varphi$ twice gives a function which is smaller than $\varphi$ in the convex partial ordering. There are some analogies present in the literature between the Fenchel conjugation operator (for which $\varphi^{**}=\varphi$) and the LLT operator, e.g. Section B.2, [CCS+22] and Section 3.1, [GLL+25]. Unfortunately, essentially any non-quadratic example witnesses $\varphi^{\sharp\sharp}\neq\varphi$. [^3] An easy way to see this is to take any $\varphi$ that is not self-concordant. Then $\varphi^{\sharp\sharp}=\varphi$ is ruled out simply by Fact 2.
@@
Now let $\gamma_{k+0.5}$ be the coupling of $(\mu_{k+0.5},\mu^{\prime}_{k+0.5})$ which first draws $(\mathbf{x},\mathbf{x}^{\prime})\sim\gamma_{k}$, and then conditioned on these values, draws and outputs $(\mathbf{y},\mathbf{y}^{\prime})$ from the coupling realizing $W^{\varphi^{*\tau}}(\tilde{\pi}_{(\tau)}^{Y\mid X=\mathbf{x}}\|\tilde{\pi}_{(\tau)}^{Y|X=\mathbf{x}^{\prime}})$.
Conditioning on the values of $(\mathbf{y},\mathbf{y}^{\prime})\sim\gamma_{k+0.5}$, we have respectively by $V+\tau\psi\succeq(1+\frac{\alpha}{\tau})\tau\psi$, Fact 13, and Fact 12 that
@@ Appendix D Comparison of Lemma 10 with Gaussian Setting
We use the second line of (29) and follow the proof of Proposition 1, [STZ25]. Here, $\tilde{\pi}_{(\tau)}^{Y}$ is almost of the form of a Gaussian convolution. First define
@@
Then this is the convolution of two densities that satisfy Poincaré inequalities with constants $\alpha$ and $\tau$, respectively. Therefore, by Lemma 2, $\nu_{\tau}$ satisfies a $\frac{\alpha\tau}{\alpha+\tau}$-Poincaré inequality. Now $\nu_{\tau}$ and $\tilde{\pi}_{(\tau)}^{Y}$ are equivalent up to a $\frac{1}{\tau}$-Lipschitz mapping; thus, by Proposition 2.3.3, [CHE23], $\tilde{\pi}_{(\tau)}^{Y}$ satisfies a $\frac{\alpha}{\tau(\alpha+\tau)}$-Poincaré inequality. ∎
@@
The proof via convexity is brittle in two respects. Generally, Assumption 2 may not hold. Moreover, it is not necessarily the case that $\nabla^{2}\varphi^{*\tau}\succeq\frac{1}{\tau}\nabla^{2}\varphi$. Indeed, consider the proof of Lemma 25, [STZ25] for general $\varphi$. We would like